Kofi and kweku are two brothers.kofi is older kweku. given that kofi's age is (5x- 4)years and kweku's age is (2x+1)years,
a.write down an expression, in terms of x,for how much older kofi is than kweku.simplify your answer
b.if kofi is ten years oder than kweku,calculate,
i.the value of x
ii.the ages of kofi and kweku

Answer by stanbon(48564)

kofi is older kweku. given that kofi's age is (5x- 4)years and kweku's age is (2x+1)years,
a.write down an expression, in terms of x,for how much older kofi is than kweku.simplify your answer
(5x-4)-(2x+1) = 3x-5
-----------------------------------------------
b.if kofi is ten years older than kweku,calculate,
i.the value of x
3x-3=10
3x=13
x=13/3
----------------------------------
ii.the ages of kofi and kweku
Kofi: 5x-4= 5(13/3)-4 = (65/3)-(12/3) = 53/3
Kweku: 2x+1 = 2(13/3)+1 = 26/3 + 3/3 = 29/3
==============================
Cheers,
Stan H.